OpenStudy (anonymous):

Divide 24b^5f^6+36b^3f^4 over 4b^3f

OpenStudy (anonymous):

\[\frac{ 24b ^{5}f ^{6}+36b ^{3}f ^{4} }{ 4b ^{3}f }\]

sam (.sam.):

\[\frac{ 24b ^{5}f ^{6}+36b ^{3}f ^{4} }{ 4b ^{3}f } \\ \\ \frac{ 24b ^{5}f ^{6}}{ 4b ^{3}f }+ \frac{36b ^{3}f ^{4} }{4b ^{3}f } \\ \\ \frac{ \cancel{24}^6b ^{\cancel5^2}f ^{\cancel6^5}}{ \cancel{4b ^{3}f} }+ \frac{\cancel{36}^9b ^{\cancel{3}^1}f ^{\cancel4^3} }{\cancel{4b ^{3}f }} \]

sam (.sam.):

Which leaves you with \[6b^2f^5+9b f^3\]

sam (.sam.):

wait should be \[6b^2f^5+9 f^3\]

OpenStudy (anonymous):

What'd you do to the exponents? Add or subtract

sam (.sam.):

b^3 and b^3 cancel out

sam (.sam.):

if divide, subtract , if multiply, add

OpenStudy (anonymous):

Oh okay thanks!

sam (.sam.):

\[a^m(a^n)=a^{m+n} \\ \\ \frac{a^m}{a^n}=a^{m-n}\]

OpenStudy (anonymous):

I'll note that thanks again